OpenStudy (fionacndg):

A regular pentagon has side length 12 cm. The perimeter of the pentagon is 60 cm and the area is 247.7 cm2. A second pentagon has side lengths that are one and a half times the first pentagon. Find the perimeter of the second pentagon.

OpenStudy (fibonaccichick666):

so, a pentagon has how many sides to start?

OpenStudy (fionacndg):

5

OpenStudy (fionacndg):

@FibonacciChick666

OpenStudy (fibonaccichick666):

and how do we calculate perimeter?

OpenStudy (anonymous):

This is simple. 60*1.5=?

OpenStudy (fionacndg):

umm a+b+c... =perimeter of a pentagon @FibonacciChick666

OpenStudy (fionacndg):

40 @OrangeMaster

OpenStudy (anonymous):

No, it would be 90

OpenStudy (fibonaccichick666):

yup, so now, we know each side is 1.5 times the old side. So just add it up

OpenStudy (fibonaccichick666):

@OrangeMaster , clearly they don't understand so why would you just give them the answer?

OpenStudy (fionacndg):

oops divided @OrangeMaster

OpenStudy (anonymous):

I didn't. I showed them how to do it. There is no need to do what you did.

OpenStudy (fionacndg):

@FibonacciChick666 13.5 each side?

OpenStudy (fibonaccichick666):

You added, try times-ing. (multiplication)

OpenStudy (anonymous):

60*1.5 is all you needed.

OpenStudy (anonymous):

Then to find the length of each side:\[90\div5=18\]

OpenStudy (fibonaccichick666):

and it won't help because they don't understand the concept. Yes you get the same answer, but concept wise is what I am going for

OpenStudy (fionacndg):

@FibonacciChick666 18

OpenStudy (anonymous):

This is the same concept, but much faster. He knows what the length of each side and the total length, but much faster.

OpenStudy (fibonaccichick666):

yes good, so because it was a regular pentagon, that means that all of the sides have the same length, so now, what do you get for the perimeter?

OpenStudy (fionacndg):

ohhh @FibonacciChick666 90

OpenStudy (anonymous):

Perimeter: 90 Individual Side Length: 18 Are we good now?

OpenStudy (fibonaccichick666):

yep, so as Orange pointed out you can do this another way: because you could factor a 1.5 out of each term, you could just multiply by 1.5

OpenStudy (anonymous):

His way was also good. I guess everyone has their own way.

OpenStudy (fionacndg):

now I get this!! thanks so much! @OrangeMaster @FibonacciChick666

OpenStudy (fionacndg):

I would have taken any help I can get

OpenStudy (fibonaccichick666):

np

OpenStudy (anonymous):

same here

OpenStudy (anonymous):

Side scale of the 2 pentagon: 3/2 Perimeter of the small pentagon: 60 Perimeter of the big one: 60 x (3/2) = 90 Area of small pentagon: 247.7 Area of big one: (247.7) x (3/2)^2 = 247.7 x (9/4) = ..... Reminder: area scale is the square of side scale.